OpenStudy (el_arrow):

find f f''(x)=2+cosx, f(0)=-1, f(pi/2)=0

OpenStudy (freckles):

can you find f' given f''

OpenStudy (el_arrow):

i got it down to the second anti-derivative x^2-cosx+Cx+D

OpenStudy (el_arrow):

@freckles

OpenStudy (fibonaccichick666):

...? not sure how you got that. Can you give your steps

OpenStudy (el_arrow):

the answer on the back of the book is x^2-cosx-1/2pix

OpenStudy (el_arrow):

i need help with getting the -1/2pix

OpenStudy (anonymous):

first step is \[f'(x)=2x+\sin(x)+C\] and find \(C\) but i think there is a typo in your question (maybe not) are you told a value of the derivative?

OpenStudy (fibonaccichick666):

ohhh, it's a double prime, it only looked like a single

OpenStudy (el_arrow):

yeah they are f(0)=-1 and f(pi/2)=0

OpenStudy (fibonaccichick666):

those are both only for f not the derivative

OpenStudy (el_arrow):

yeah just f

OpenStudy (anonymous):

ok then we have more work to do

OpenStudy (anonymous):

\[f(x)=x^2-\cos(x)+Cx+D\] and you need \(C\) and \(D\)

OpenStudy (el_arrow):

yeah

OpenStudy (anonymous):

D is easy enough to find, since you know \[f(0)=-1\]

OpenStudy (el_arrow):

yeah i got that one

OpenStudy (anonymous):

i get \(D=0\) do you?

OpenStudy (el_arrow):

no i got d=-1

OpenStudy (anonymous):

ok \[f(0)=0^2-\cos(0)+C\times 0+D=-1\]

OpenStudy (anonymous):

since \(\cos(0)=1\) you have \[-1+D=-1\]

OpenStudy (el_arrow):

oh no my mistake

OpenStudy (el_arrow):

so d=0 what about c

OpenStudy (anonymous):

now we have \[f(x)=x^2-\sin(x)+Cx\] and you can find \(D\) by using the fact that \[f(\frac{\pi}{2})=0\]

OpenStudy (anonymous):

i meant of course "find C"

OpenStudy (el_arrow):

i get c=pi/4

OpenStudy (el_arrow):

but the answer is x^2-cosx-1/2pix

OpenStudy (anonymous):

ok that was wrong

OpenStudy (el_arrow):

but cos of pi/2 is 0

OpenStudy (anonymous):

damn another typo \[f(\frac{\pi}{2})=\frac{\pi^2}{4}+\frac{\pi}{2}C=0\]

OpenStudy (el_arrow):

lol

OpenStudy (anonymous):

that makes \[\frac{\pi}{2}C=-\frac{\pi^2}{4}\]

OpenStudy (anonymous):

making \(C=-\frac{\pi}{2}\)